OpenStudy (anonymous):

Please help me with these im not good at these at all :P 1. If a vector has an initial point (0,0) and a terminal point (-15,8) which is closest to the magnitude of the vector? A. 7.0 B. 12.7 C. 17.0 D.23.0 2. write the sum of the two vectors as an ordered pair. (7,-5) and (4,4) A. (-1,11) B. (3,-1) C. (11,-1) D. (-1,3)

hero (hero):

f you have two points P = (p_1, p_2) and Q = (q_1, q_2), of line segment PQ, then vector v is represented by <v1, v2> = <q1 - p1, q2 - p2> and the length of vector v is: \[||v|| = \sqrt{(q_1 - p_1)^2 + (q_2 - p_2)^2}\] \(= \sqrt{v^2_1 + v^2_2}\)

hero (hero):

In this case P = (0,0) and Q = (-15,8) so \[||v|| = \sqrt{(-15 - 0)^2 + (8 - 0)^2}\]

OpenStudy (anonymous):

Its 17 right?